OpenStudy (anonymous):

when u are having in this form then how to differentiate :) tan^-1(2^(x+1)/(1-4^x)) can i use substitution here ?

OpenStudy (anonymous):

i don't think so substitution will work

hartnn (hartnn):

substitute 2^x = tan y

hartnn (hartnn):

lol, why not ?

hartnn (hartnn):

do you know tan 2 theta formula ?

OpenStudy (anonymous):

coz i have never did this type of question lol

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

2tanα/(1 – tan²α)

hartnn (hartnn):

2^(x+1) = 2. 2^x , 4^x = (2^x)^2 plug in 2^x = tan y and see the magic! (simplification :P)

OpenStudy (anonymous):

will it really works :PPPP lol i hate chain rule -_-

OpenStudy (anonymous):

don't wanna use chain rule for inverse trig ...........

hartnn (hartnn):

? what did you get after simplification ?

hartnn (hartnn):

i don't think chain rule is required here...

OpenStudy (anonymous):

wait :))

OpenStudy (anonymous):

use tan(x+y)=(tanx+tany)/(1-tanxtany)

OpenStudy (anonymous):

@Abhishek619 :/

hartnn (hartnn):

same thing....tan 2 thetha comes from tan (x+y) formula only

OpenStudy (anonymous):